OpenStudy (ajanijones):

help i give medals asap... What are the diameter and radius of a circle with a circumference of 32π ft? A. d = 16 ft r = 8 ft B. d = 32 ft r = 16 ft C. d = 16 ft r = 32 ft D. d = 8 ft r = 16 ft

OpenStudy (anonymous):

C = pi * d or 2*pi*r 32 pi = pi* d 32 = d d/2 = r 32/2 = 16 16 = r

OpenStudy (ajanijones):

Ok yea i though soo... thx i gave you medal.